5 mm mm 0.5 mm 1.- 3.04 2.0 100 x X X

calculate the magnification of fields 1 and 3, and the field diameter of 2 . (hint. Use your ruler.) Note the the numbers for the field diameters below are too large to represent the typical compound microscope lens system, but the relationships depicted are accurate
Answer #1

For a microscope

m\times d=constant..........(1)

With m = magnification of microscope

d= diameter of field of view

Given for field 1

\\d_1=5mm\\\ m_1=?

For field 2

\\d_2=?\\\ m_2=100

For field 3

\\d_3=0.5mm\\ m_3=?

therefore from equation (1)

\\m_1d_1=m_2d_2\\and \\ m_2d_2=m_3d_3

Put the values from above we get

\\m_1\times 5=100\times d_2\\and \\ 100\times d_2=m_3\times 0.5

Or

\\m_1=20d_2\\and \\ m_3=200d_2

Now from figure we can see

5mm >d_2>0.5mm

So let's choose d_2=2.5mm

Therefore From above equation

\\m_1=20\times 2.5=50\\and \\ m_3=200\times 2.5=500

Therefore for field diameter=2.5mm of #2

Magnification for field #1 and #3 are 50 and 500 respectively.

Similarly you can choose other possible value of d_2 as per your ruler

And correspondingly magnification of field 1 and 3.